OpenStudy (anonymous):

integral of 1/(1-cosx)????

OpenStudy (anonymous):

multiply top and botom by conjugate (1-cosx) int(1-cosx)/sin^2xdx =intcsc^2x-intcosx/sin^2xdx for 2nd part use sub u=sinx du=cosxdx sub intcosx/sin^2xdx=intdu/u^2 =-1/u =-1/sinx=-cscx the answer is -cotx-cscx+c

OpenStudy (anonymous):

Alternatively, you can use the half-angle substitution \(t=\tan\dfrac{x}{2}\) to determine an expression for \(\cos x\) in terms of \(t\) and end up with a rational function of \(t\).

OpenStudy (anonymous):

@Lilium I believe you meant \(1+\cos x\) as the conjugate.